A 10.0 x10^(-6)F capacitor is charged by a 10.0V battery through a resistance R. The capacitor reaches a potential difference of 4.00V in a time interval of 3.00s after charging begins. Find R.
1
Expert's answer
2021-02-22T10:29:38-0500

The voltage across the capacitor at any instant in time during the charging period can be found as follows:

"V_c=V_s(1-e^{-\dfrac{t}{RC}}),""4=10\cdot(1-e^{-\dfrac{3}{RC}}),""e^{-\dfrac{3}{RC}}=0.6,""ln(e^{-\dfrac{3}{RC}})=ln(0.6),""-\dfrac{3}{RC}=ln(0.6),""R=\dfrac{-3}{ln(0.6)\cdot10\cdot10^{-6}\ F}=587.28\ k\Omega."